OpenStudy (anonymous):

How many hydrogen atoms are present in a 22.4 gram sample of propane, C3H8?

OpenStudy (abb0t):

First, you need to calculate the number of mol's present in hydrogen. Then, use avogadro's number to find the number of atoms. This is how you will solve it: 22.4 g of Propane \(\times\) \(\frac{moles}{mass~of~propane}\)\(\times\)\(\frac{6.022 \times 10^{23} atoms}{mole}\)